EINSTEIN A-COEFFICIENTS AND STATISTICAL WEIGHTS FOR MOLECULAR ABSORPTION TRAN-SITIONS IN THE HIT RAN DATABASE

\underline{MARIE SIMECKOVA}, LAURENCE S. ROTHMAN, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA; DAVID JACQUEMART, Laboratoire de Dynamiques, Interactions et Réactivités, Université Pierre at Marie Curie, 4 place Jussieu, Paris 75252, France; ROBERT R. GAMACHE, University of Mass. Lowell, Department of Environmental, Earth & Atmospheric Sciences, Lowell MA 01854, USA; AARON GOLDMAN, University of Denver, Department of Physics, Denver CO 80208, USA.

The weighted square of the transition moment presented in previous editions of HIT RAN has been replaced by the Einstein A-coefficient in HIT RAN 2004. The calculation has been performed using the HIT RAN line intensities S_{HIT}, the statistical weights g_{2} of the upper levels and other related quantities

A_{21} = 8πcν_{0}^{2}Q_{tot}(T_{0})S_{HIT} /(e^{−c_{2}E_{1}/T_{0}}(1 − e^{−c_{2}ν_{0}/T_{0}})I_{a}g_{2}).

The motivations and the corresponding theory, including the determination of the statistical weights for molecules of various symmetries, will be discussed.